# Multi-qubit Dynamical Decoupling for Enhanced Crosstalk Suppression

### 08 Mar 2024 v1

@@ Abstract
Dynamical decoupling (DD) is one of the simplest error suppression methods, aiming to enhance the coherence of qubits in open quantum systems. Moreover, DD has demonstrated effectiveness in reducing coherent crosstalk, one major error source in near-term quantum hardware, which manifests from two types of interactions. Static crosstalk exists in various hardware platforms, including superconductor and semiconductor qubits, by virtue of always-on qubit-qubit coupling. Additionally, driven crosstalk may occur as an unwanted drive term due to leakage from driven gates on other qubits. Here we explore a novel staggered DD protocol tailored for multi-qubit systems that suppresses the decoherence error and both types of coherent crosstalk. We develop two experimental setups – an “idle-idle” experiment in which two pairs of qubits undergo free evolution simultaneously and a “driven-idle” experiment in which one pair is continuously driven during the free evolution of the other pair. These experiments are performed on an IBM Quantum superconducting processor and demonstrate the significant impact of the staggered DD protocol in suppressing both types of coherent crosstalk. When compared to the standard DD sequences from state-of-the-art methodologies with the application of X2 sequences, our staggered DD protocol enhances circuit fidelity by 16.9% and 8.5%, respectively, in addressing these two crosstalk types.
@@ I Introduction
The inherent noise in near-term quantum computers hinders the demonstration of practical quantum advantage. Recently, a variety of quantum error mitigation techniques have been developed to tackle this challenge, such as zero noise extrapolation (ZNE) [1, 2], probabilistic error cancellation (PEC) [3], and clifford data regression (CDR) [4], which require substantial overhead in quantum resources (such as the number of circuits executed) in order to model the noise and mitigate the resulting error. Quantum error suppression techniques, such as randomized compiling [5], pulse-scaling for Trotterized circuits [6], and dynamical decoupling (DD) [7, 8, 9], offer a reduction in error with no additional quantum resources. DD aims to reduce the decoherence error by decoupling the interaction between the quantum system and the environment. This process involves inserting a sequence of gates that compose the identity during the idle period when the qubit is not being used for computation. This entails very little classical overhead, as it only requires identifying the idle times and inserting sequences with a pattern of delays between the gates. It makes DD one of the simplest error suppression methods. Numerous DD sequences have been introduced over decades [10, 8, 11, 12, 13], leading to performance improvements on various modalities of quantum computers, including superconducting circuits [14, 15, 16, 17, 18], trapped ion [19, 20], and neutral atoms [21]. However, these studies typically focus on a single qubit coupled to its environment (with the exception of Ref. [22]), however quantum computers consist of many qubits coupled to each other. Regarding Ref. [22], the authors studied static $ZZ$ crosstalk to propose a crosstalk-robust dynamical decoupling (CRDD) protocol, similar in pattern those we consider, and demonstrate improved multi-qubit state preservation over standard DD sequences.
@@
Here, we investigate a different multi-qubit DD strategy called staggered DD that consists of a standard DD sequence with symmetric free evolution times on one qubit with the same sequence on an adjacent qubit interleaved in time such that the gates occur in the middle of the free evolution times of the original sequence. In addition to suppressing dephasing and single-qubit coherent errors, we show that staggered DD can suppress undesired two-qubit interactions by inverting their direction of rotation in the two-qubit Hilbert space during each of the half-idle times of the gate interleaving. We first provide some theoretical insight for this method, by placing it in context with the usual treatment of DD for general decoherence. We show that this can suppress crosstalk from unwanted $ZZ$ interactions, which can arise both from static coupling between transmon superconducting qubits and driven crosstalk from cross resonance interaction. We then conduct experiments on an IBM quantum computer to explore both static and driven crosstalk. Even though the standard DD sequence has already shown to reduce the decoherence error and crosstalk [18, 23], the experimental results demonstrate that staggered DD performs better than standard (unstaggered) sequences applied naïvely to all qubits.
@@ II Theoretical analysis
If instead, only a single gate were applied to each qubit at a time, the crosstalk can be inverted,
@@
The goal of the staggered X2 sequence, in which the DD sequences differ for two qubits $(q_{0},q_{1})$, is to cancel the accumulated two-qubit $ZZ$ rotation by staggering the timing of the gates in the DD sequences. Specifically, the X2 staggered sequence
@@
in which the accumulated phases cancel because each $\theta_{i}$ is an equal rotation over the free evolution time $\tau/2$. An illustration of applying the standard or staggered X2 sequence to a two-qubit circuit is shown in Fig. 1. We can also invert the staggered DD sequences between the control and target qubits to the same effect. This staggered technique can likewise be extended to other DD sequences, such as
@@
where explicit references to the free evolutions $f_{\tau}$ have been omitted for simplicity. Noisy simulations illustrating the cancellation of $ZZ$ crosstalk via the use of staggered DD is shown in Appendix C.
@@ III.1 Experimental Protocol
We select a subset of nine qubits from the 27-qubit ibm_cairo backend, which includes controlled-NOT (CX) pairs generated by both echoed cross resonance (ECR) [31, 32] and direct CX (DCX) [26]. We focus on physically connected qubit pairs due to their increased susceptibility to crosstalk. In this configuration, each pair neighbors the other, where operations on one pair can induce crosstalk in the adjacent pair [33, 34]. We present the RB experiments in two experimental settings: (i) idle-idle in which the two pairs of qubits are simultaneously idle during their simultaneous RB sequences (aside from the relevant DD sequences) in Section III.2, and (ii) idle-driven in which the fidelity of the vulnerable pair of qubits is measured after the RB sequence (with varying DD sequence) while the other pair of qubits is driven by a random sequence of two-qubit gates (SU(4)) in Section III.3. The two-qubit individual RB circuit sets a baseline for the two experiments for comparison, where the $ZZ$ interaction between qubits in the pair is suppressed by calibrating rotary echo pulses on the target qubits [35]. This is compared to the same RB circuit but while a neighboring is active, effectively exposing the pair to static and/or driven crosstalk (the effect of static crosstalk in which the neighboring qubits are in the ground state has been calibrated out by the backend). The idle-idle and idle-driven experiments expand to involve two qubit pairs, totaling four qubits. Standard and staggered DD sequences are inserted during the idle periods. Specifically, in the idle-idle experiment, DD are applied to both qubit pairs, whereas in the second experiment, DD is only applied to the first qubit pair with RB sequences. These experiments highlight the difference in performance between standard and staggered DD sequences when the qubits are exposed to both stochastic dephasing and coherent crosstalk arising from static $ZZ$ coupling and cross-resonance driven $ZZ$. The quantum circuits were generated and executed using the Qiskit Experiments framework [36].
@@
We choose a pair of qubits $(q_{11},q_{14})$ and $(q_{12},q_{13})$ for study, that are particularly vulnerable to crosstalk from operations on neighboring qubit pair, as shown by the fidelity differences between Fig. 3a and Fig. 3b. For idle-idle experiment, we apply exactly the same RB circuit simultaneously to the adjacent pair of qubits. While the duration of the idle period remains the same for both qubit pairs, the starting points of these idle periods differ slightly due to the differences in duration of the two neighboring CX pairs. We use the As Late As Possible policy to schedule the circuits. For the idle-driven experiment, while one qubit pair undergoes the RB circuit, its adjacent pair is subjected to a circuit composed of different random SU(4) blocks, which continuously drives crosstalk during the RB experiment, as shown in Fig. 2e. The number of SU(4)s in the circuit is determined to make sure that has the same duration as the RB circuit. This experimental arrangement ensures the simultaneous execution of both circuits. Since no idle time is inserted between SU(4)s, during the idle periods of the RB circuit, the neighboring qubits in the SU(4) circuit remain active thereby contributing driven crosstalk.
@@
During the idle period, DD sequences implemented in different ways at the pulse and timing level are inserted as depicted in Fig. 2b-d. Standard DD sequences have symmetric timing on both qubits, and they occur simultaneously (see Fig. 2b for a representation of the X2 sequence). For staggered sequences, the sequence on one qubit is symmetric (as in the standard case) and the other is offset such that each staggered pulse occurs halfway between the symmetric pulses on the other qubit (Fig. 2c). The “inverse” staggered sequence switches the role of symmetric and staggered sequences on the qubits (Fig. 2d). While each sequence is nominally identical from the perspective of unitary operations (composing to the identity), there are control variations at the pulse level [17] that may effect their performance. We consider the default X gate to be a $+\pi$ rotation around the $x$-axis of the Bloch sphere (also denoted Xp) and Xm a $-\pi$ rotation with negative amplitude at the pulse level. These variations consisting of alternating rotations are denoted with pm after the sequence name. For example, the X2pm sequence consists of the gates Xp-Xm.
In the two experiments, the CX gate is aligned with the native direction of either ECR or DCX gates, determined in calibration to that of higher fidelity. For the two adjacent qubit pairs, the neighbor qubits from each pair can be control-control, control-target, or targt-target in their respective pairs. The performances across these different configurations, in terms of the type of CX gate and the type of connected neighbor qubits, are detailed in Appendix A.
@@ III.2 Idle-idle Experiment - Simultaneous Randomized Benchmarking
We evaluate the performances of standard DD sequences, including X2/X2pm, XY4/XY4pm, XY8/XY8pm, and their staggered and inverse of staggered variants - 16 sequences in total. We then compare these sequences to a baseline, in which the qubits remain free evolution during the delay, with no DD sequence inserted as shown in Fig. 2a.
@@
The RB and simultaneous RB (SRB) results for the qubit pair ($q_{11},q_{14}$) are illustrated in Fig. 3, showcasing the comparison between various X2 sequences (X2pm, X2pm-stag, X2pm-stag-inv) against the free evolution. In the SRB experiment, it involves the adjacent qubit pair of ($q_{11},q_{14}$), which is ($q_{12}$, $q_{13}$), with $q_{14}$ and $q_{13}$ being the target qubits in their respective pairs and neighboring each other. The figure shows the connectivity of the two qubit pairs, detailing the type of CX gate used (ECR or DCX) and identifying whether a qubit acts as a control (C) or target (T) for the neighboring qubit connections between pairs. The connection between the qubit pair ($q_{11},q_{14}$) is highlighted in red to indicate the specific results corresponding to this pair in the figure. The insertion of default X2 pulses results in unexpected behavior, even degrading the performance below that of the free evolution. The reason for this outcome is elaborated in Appendix B.
@@
We also demonstrate the results for XY4 and XY8 sequences in Fig. 4. Given the negligible difference between the results for staggered DD and its inverse, only the staggered DD is depicted. We observe “dropouts” in fidelity for a short idle times in the XY8 sequence which can be explained by constructive interference of higher frequency spectral noise components [37, 38]. In addition, the XY8 sequence shows consistently worse results than XY4.. Therefore, our subsequent discussions will focus primarily on the X2 and XY4 sequences. The correlation between the duration of the idle period and various DD sequences is elaborated upon in Appendix LABEL:app:order.
@@
Figure 3: Free evolution experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with different DD sequences: X2pm, X2pm staggered, and X2pm staggered inverse. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs. The significant difference in fidelity during free evolution between the RB and SRB experiments indicates additional crosstalk generated by the neighboring qubit pair.
@@
Figure 4: Free evolution experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with XY4pm, XY4pm staggered, XY8pm, and XY8pm staggered. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs.
III.3 Driven-idle Experiment - Randomized Benchmarking and Random SU(4
@@
Figure 5: Crosstalk experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when only RB is applied. (b) The same probability when of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair while SU(4) is applied to its adjacent qubit pair $(q_{12},q_{13})$.
@@
In Fig. 5, we present the fidelity results when (1) applying only the RB circuit to ($q_{11},q_{14}$) and (2) additionally performing the SU(4) circuit on the adjacent ($q_{12},q_{13}$) pair. Again, the qubits neighboring each other across the two pairs are target qubits within their respective pairs. Similar to the free evolution experiment, the staggered DD sequences consistently outperform the non-staggered DD sequences. In comparison to the baseline, we observe a significant enhancement in fidelity for the X2pm, XY4pm, X2pm-stag, and XY4pm-stag sequences when the RB circuit operates independently. This improvement becomes more pronounced when the RB and SU(4) circuits are executed simultaneously.

### 14 Mar 2024 v2

@@ Abstract
Dynamical decoupling (DD) is one of the simplest error suppression methods, aiming to enhance the coherence of qubits in open quantum systems. Moreover, DD has demonstrated effectiveness in reducing coherent crosstalk, one major error source in near-term quantum hardware, which manifests from two types of interactions. Static crosstalk exists in various hardware platforms, including superconductor and semiconductor qubits, by virtue of always-on qubit-qubit coupling. Additionally, driven crosstalk may occur as an unwanted drive term due to leakage from driven gates on other qubits. Here we explore a novel staggered DD protocol tailored for multi-qubit systems that suppresses the decoherence error and both types of coherent crosstalk. We develop two experimental setups – an “idle-idle” experiment in which two pairs of qubits undergo free evolution simultaneously and a “driven-idle” experiment in which one pair is continuously driven during the free evolution of the other pair. These experiments are performed on an IBM Quantum superconducting processor and demonstrate the significant impact of the staggered DD protocol in suppressing both types of coherent crosstalk. When compared to the standard DD sequences from state-of-the-art methodologies with the application of X2 sequences, our staggered DD protocol enhances circuit fidelity by 19.7% and 8.5%, respectively, in addressing these two crosstalk types.
@@ I Introduction
The inherent noise in near-term quantum computers hinders the demonstration of practical quantum advantage. Recently, a variety of quantum error mitigation techniques have been developed to tackle this challenge, such as zero noise extrapolation (ZNE) [1, 2], probabilistic error cancellation (PEC) [3], and clifford data regression (CDR) [4], which require substantial overhead in quantum resources (such as the number of experiments executed) in order to model the noise and mitigate the resulting error. Quantum error suppression techniques, such as randomized compiling [5], pulse-scaling for Trotterized circuits [6], and dynamical decoupling (DD) [7, 8, 9], offer a reduction in error with no additional quantum resources. DD aims to reduce the decoherence error by decoupling the interaction between the quantum system and the environment. This process involves inserting a sequence of gates that compose the identity during the idle period when the qubit is not being used for computation. This entails very little classical overhead, as it only requires identifying the idle times and inserting sequences with a pattern of delays between the gates. It makes DD one of the simplest error suppression methods. Numerous DD sequences have been introduced over decades [10, 8, 11, 12, 13], leading to performance improvements on various modalities of quantum computers, including superconducting circuits [14, 15, 16, 17, 18, 19], trapped ions [20, 21], and neutral atoms [22]. These studies typically focus on a single qubit and its environment but quantum computers consist of many qubits coupled to each other. Notable exceptions [23, 18] explore static $ZZ$ crosstalk in multi-qubit systems. They propose dynamical decoupling (DD) protocols to protect the quantum systems from crosstalk and enhance the gate fidelity, similar in pattern to those we consider.
@@
Here, we investigate a different multi-qubit DD strategy called staggered DD that consists of a standard DD sequence with symmetric free evolution times on one qubit with the same sequence on an adjacent qubit interleaved in time such that the gates occur in the middle of the free evolution times of the original sequence. In addition to suppressing dephasing and single-qubit coherent errors, we show that staggered DD can suppress undesired two-qubit interactions by inverting their direction of rotation in the two-qubit Hilbert space during each of the half-idle times of the gate interleaving. We first provide some theoretical insight for this method, by placing it in context with the usual treatment of DD for general decoherence. We show that this can suppress crosstalk from unwanted $ZZ$ interactions, which can arise both from static coupling between transmon superconducting qubits and driven crosstalk from cross resonance interaction. We then conduct experiments on an IBM quantum computer to explore both static and driven crosstalk. Even though the standard DD sequence has already shown to reduce the decoherence error and crosstalk, the experimental results demonstrate that staggered DD performs better than standard (unstaggered) sequences applied naïvely to all qubits.
@@ II Theoretical analysis
If instead, only a single gate is applied to each qubit at a time, the crosstalk can be inverted,
@@
The goal of the staggered X2 sequence, in which the DD sequences differ for two qubits $(q_{0},q_{1})$, is to cancel the accumulated two-qubit $ZZ$ rotation by staggering the timing of the gates in the DD sequences. Specifically, the staggered X2 sequence
@@
in which the accumulated phases cancel because each $\theta_{i}$ is an equal rotation over the free evolution time $\tau/2$. An illustration of applying the standard or staggered X2 sequence to a two-qubit circuit is shown in figure 1. We can also invert the staggered DD sequences between the control and target qubits to the same effect. This staggered technique can likewise be extended to other DD sequences, such as
@@
where explicit references to the free evolutions $f_{\tau}$ have been omitted for simplicity. Noisy simulations illustrating the cancellation of $ZZ$ crosstalk via the use of staggered DD are shown in Appendix C.
@@ III.1 Experimental Protocol
We select a subset of nine qubits from the 27-qubit ibm_cairo backend, which includes controlled-NOT (CX) pairs generated by both echoed cross resonance (ECR) [31, 32] and direct CX (DCX) [26]. We focus on physically connected qubit pairs due to their increased susceptibility to crosstalk. In this configuration, each pair neighbors the other, where operations on one pair can induce crosstalk in the adjacent pair [33, 34, 35]. We present the RB experiments in two experimental settings: (i) idle-idle in which the two pairs of qubits are simultaneously idle during their simultaneous RB sequences (aside from the relevant DD sequences) in Section III.2, and (ii) idle-driven in which the fidelity of the vulnerable pair of qubits is measured after the RB sequence (with varying DD sequence) while the other pair of qubits is driven by a random sequence of two-qubit gates (SU(4)) in Section III.3. The two-qubit individual RB circuit sets a baseline for the two experiments for comparison, where the $ZZ$ interaction between qubits in the pair is suppressed by calibrating rotary echo pulses on the target qubits [36]. This is compared to the same RB circuit but while a neighboring qubit pair is active, effectively exposing the RB circuit to static and/or driven crosstalk (the effect of static crosstalk in which the neighboring qubits are in the ground state has been calibrated out by the backend). The idle-idle and idle-driven experiments expand to involve two qubit pairs, totaling four qubits. Standard and staggered DD sequences are inserted during the idle periods. Specifically, in the idle-idle experiment, DD are applied to both qubit pairs, whereas in the second experiment, DD is only applied to the first qubit pair with RB sequences. These experiments highlight the difference in performance between standard and staggered DD sequences when the qubits are exposed to both stochastic dephasing and coherent crosstalk arising from static $ZZ$ coupling and cross-resonance driven $ZZ$. The quantum circuits were generated and executed using the Qiskit Experiments framework [37].
@@
We choose a pair of qubits $(q_{11},q_{14})$ and $(q_{12},q_{13})$ for study, that are particularly vulnerable to crosstalk from operations on neighboring qubit pair, as shown by the fidelity differences between figure 3a and figure 3b. For idle-idle experiment, we apply exactly the same RB circuit simultaneously to the adjacent pair of qubits. While the duration of the idle period remains the same for both qubit pairs, the starting points of these idle periods differ slightly due to the differences in duration of the two neighboring CX pairs. We use the As Late As Possible policy to schedule the circuits. For the idle-driven experiment, while one qubit pair undergoes the RB circuit, its adjacent pair is subjected to a circuit composed of different random SU(4) blocks, which continuously drives crosstalk during the RB experiment, as shown in figure 2e. The number of SU(4)s in the circuit is determined to ensure it has the same duration as the RB circuit. This experimental arrangement ensures the simultaneous execution of both circuits. Since no idle time is inserted between SU(4)s, during the idle periods of the RB circuit, the neighboring qubits in the SU(4) circuit remain active thereby contributing to driven crosstalk.
@@
During the idle period, DD sequences implemented in different ways at the pulse and timing level are inserted as depicted in figure 2b-d. Standard DD sequences have symmetric timing on both qubits, and they occur simultaneously (see figure 2b for a representation of the X2pm sequence). For staggered sequences, the sequence on one qubit is symmetric (as in the standard case) and the other is offset such that each staggered pulse occurs halfway between the symmetric pulses on the other qubit (figure 2c). The “inverse” staggered sequence switches the role of symmetric and staggered sequences on the qubits (figure 2d). While each sequence is nominally identical from the perspective of unitary operations (composing to the identity), there are control variations at the pulse level [17] that may affect their performance. We consider the default X gate to be a $+\pi$ rotation around the $x$-axis of the Bloch sphere (also denoted Xp) and Xm a $-\pi$ rotation with negative amplitude at the pulse level. These variations consisting of alternating rotations are denoted with pm after the sequence name. For example, the X2pm sequence consists of the gates Xp-Xm.
In the two experiments, the CX gate is aligned with the native direction of either ECR or DCX gates, determined in calibration to that of higher fidelity. For the two adjacent qubit pairs, the neighbor qubits from each pair can be control-control, control-target, or target-target in their respective pairs. The performances across these different configurations, in terms of the type of CX gate and the type of connected neighbor qubits, are detailed in Appendix A.
@@ III.2 Idle-idle Experiment - Simultaneous Randomized Benchmarking
We evaluate the performances of standard DD sequences, including X2/X2pm, XY4/XY4pm, XY8/XY8pm, and their staggered and inverse of staggered variants - 18 sequences in total. We then compare these sequences to a baseline, in which the qubits remain free evolution during the delay, with no DD sequence inserted as shown in figure 2a.
@@
The RB and simultaneous RB (SRB) results for the qubit pair ($q_{11},q_{14}$) are illustrated in figure 3, showcasing the comparison between various X2 sequences (X2pm, X2pm-stag, X2pm-stag-inv) against the free evolution. In the SRB experiment, it involves the adjacent qubit pair of ($q_{11},q_{14}$), which is ($q_{12}$, $q_{13}$), with $q_{14}$ and $q_{13}$ being the target qubits in their respective pairs and neighboring each other. The figure shows the connectivity of the two qubit pairs, detailing the type of CX gate used (ECR or DCX) and identifying whether a qubit acts as a control (C) or target (T) for the neighboring qubit connections between pairs. The connection between the qubit pair ($q_{11},q_{14}$) is highlighted in red to indicate the specific results corresponding to this pair in the figure. The insertion of default X2 pulses results in unexpected behavior, even degrading the performance below that of the free evolution. The reason for this outcome is elaborated in Appendix  B.
@@
We also demonstrate the results for XY4 and XY8 sequences in figure 4. Given the negligible difference between the results for staggered DD and its inverse, only the staggered DD is depicted. We observe “dropouts” in fidelity for a short idle times in the XY8 sequence which can be explained by constructive interference of higher frequency spectral noise components [38, 39]. In addition, the XY8 sequence shows consistently worse results than XY4. Therefore, our subsequent discussions will focus primarily on the X2 and XY4 sequences.
@@
Figure 3: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with different DD sequences: X2pm, staggered X2pm, and inverse of staggered X2pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs. The significant difference in fidelity during free evolution between the RB and SRB experiments indicates additional crosstalk generated by the neighboring qubit pair.
@@
Figure 4: Idle-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$ when comparing the baseline (free evolution) with XY4pm, staggered XY4pm, XY8pm, and staggered XY8pm. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair in isolation. (b) The same probability when SRB is applied to both qubit pairs.
III.3 Driven-idle Experiment - Randomized Benchmarking and Random SU(4)
@@
Figure 5: Driven-idle experiment results on qubit pairs $(q_{11},q_{14})$ and $(q_{12},q_{13})$. (a) The probability of yielding $|00\rangle$ for $(q_{11},q_{14})$ when only RB is applied. (b) The same probability when of yielding $|00\rangle$ for $(q_{11},q_{14})$ when RB is applied to this qubit pair while SU(4) is applied to its adjacent qubit pair $(q_{12},q_{13})$.
@@
In figure 5, we present the fidelity results when (1) applying only the RB circuit to ($q_{11},q_{14}$) and (2) additionally performing the SU(4) circuit on the adjacent pair ($q_{12},q_{13}$). Again, the qubits neighboring each other across the two pairs are target qubits within their respective pairs. Similar to the idle-idle experiment, the staggered DD sequences consistently outperform the non-staggered DD sequences. In comparison to the baseline, we observe a significant enhancement in fidelity for the X2pm, XY4pm, X2pm-stag, and XY4pm-stag sequences when the RB circuit operates independently. This improvement becomes more pronounced when the RB and SU(4) circuits are executed simultaneously.